# Riemannian Time Warping: Multiple Sequence Alignment in Curved Spaces

### v1

@@ I Introduction
Time warping describes the temporal alignment of several signals – a fundamental operation for processing sequential data commonly used in speech recognition [1, 2], human action recognition [3, 4], or robot motion learning [5, 6]. Given multiple signals with varying lengths describing the same underlying process, each signal is warped with the goal to improve their temporal alignment. Identifying these warping parameters belongs to the domain of unsupervised learning.
@@
Dynamic Time Warping (DTW) [7], introduced in 1978, finds the optimal alignment analytically through Dynamic Programming [8], and is a prominent approach when dealing with just two signals, however, its computational complexity scales exponentially, which prevents its application to larger datasets. For this reason, alternative algorithms have been proposed that sacrifice global optimality to accelerate the search process by several orders of magnitude [9, 10, 11].
@@
Almost all related works focus on data in the Euclidean space. However, many real-world problems describe temporal phenomena evolving on Riemannian manifolds (cf. Figure LABEL:fig:overviewRTW). This is especially important in robotics, where many fundamental problems involve non-Euclidean geometry [12]. Typical examples in robotics are the unit sphere $\mathbb{S}^{3}$ for representing orientations or symmetric positive definite matrices $\mathbb{S}^{D}_{++}$ for representing stiffness gains or manipulability ellipsoids. Taking into account the data structure is important in data-driven machine learning and can be seen as prior geometric knowledge. In this line, Quaternion Dynamic Time Warping (QDTW) [13] extends its predecessor DTW to signals of unit quaternions, but suffers from the same computational complexity. To the best of our knowledge, so far no generic time warping method has been reported that can handle structured non-Euclidean data in general. One reason for this could be that the sum and scalar multiplication operations used in related works are not defined on Riemannian manifolds [14].
As our main contribution, we present a novel approach called Riemannian Time Warping (RTW) for efficient temporal alignment of multiple signals with varying lengths. Our method, for the first time, explicitly considers the Riemannian manifold ${\mathcal{M}}$ on which the data resides. This letter presents an extensive evaluation on challenging benchmarks, involving datasets in $\mathbb{R}^{1}$, $\mathbb{R}^{2}$, $\mathbb{S}^{1}$, $\mathbb{S}^{3}$, and $\mathbb{S}^{2}_{++}$, as well as problems with multiple signals capturing complex patterns. In averaging and classification tasks, RTW consistently outperforms the state-of-the-art [10, 11] w.r.t. alignment quality and mean signal representation. Furthermore, RTW achieves near-optimal results compared to techniques based on Dynamic Programming such as [13], however, with significantly reduced computation time, as its computational complexity scales linearly with the number of signals and data points. We also show that learning movement primitives benefits from data-preprocessing with RTW. Finally, we validate our approach in the context of real-robot motion learning with a collaborative torque-controlled KUKA LBR iiwa manipulator, leveraging the Riemannian manifold associated with rigid body motions in $\mathrm{SE}(3)$. For conducting all of these comparisons, we also extended state-of-the-art methods. An accompanying video of this work is available at https://youtu.be/v_k2LwEDH9I.
@@
Notation: In the remainder, we will use square bracket notation to describe discrete values, e.g, ${\mathbf{x}}[t]\in\mathbb{R}^{D},{\bm{x}}[t]\in{\mathcal{M}}$, and round bracket notation for continuous functions, e.g, ${\mathbf{\phi}}(z)$.
@@ II Related Work
Many related works focus on aligning ${N}=2$ signals of length ${T}$. Despite its age, the DTW approach [7] remains relevant and is widely used due to its simplicity and effectiveness. Extending DTW to ${N}>2$ signals introduces significant computational challenges, i.e. refer to Multiple Multi-Dimensional Dynamic Time Warping (MMDDTW) [15], which computes the distance matrix over all ${N}$ dimensions. While this results in a globally optimal solution through Dynamic Programming, it becomes unfeasible for $N\gg 2$ due to its computational complexity of $\mathcal{O}({T}^{{N}})$. The Non-Linear Alignment and Averaging Framework (NLAAF) [16] is another extension that iteratively aligns pairs of signals while adjusting previous solutions, with a complexity of $\mathcal{O}({N}{T}^{2})$.
Generalized Time Warping (GTW) [9], Trainable Time Warping (TTW) [10], and Neural Time Warping (NTW) [11] are newer methods that align many signals in linear time. GTW uses a pre-defined set of basis functions, which is cumbersome for complex warpings as many basis functions must be defined beforehand. TTW and NTW, on the other hand, first transform the discrete signals into the continuous time domain through sinc interpolation, apply an index shift in the continuous space, and finally sample the shifted signal back into the discrete time domain. TTW models the warpings by estimating coefficients of the discrete sine transform, while NTW uses a neural network. Both algorithms outperform GTW [10, 11], therefore we do not consider GTW in our experiments.
@@ III-A Problem Statement
Consider ${N}$ discrete signals ${\mathbf{X}}=\{{\mathbf{x}_{n}}\}_{n=1}^{N}$ of lengths ${T_{n}}$, i.e. ${\mathbf{x}_{n}}=\{{\mathbf{x}_{n}}[t]\}_{t=1}^{{T_{n}}}$, where ${\mathbf{x}_{n}}[t]\in\mathbb{R}^{D}$. Signals with temporal modifications ${\mathbf{{\hat{X}}}}=\{{\mathbf{{\hat{x}}}_{n}}\}_{n=1}^{N}$ of new length ${Z}\geq{T_{\text{max}}}=\max{\left\{{T_{n}}\right\}}_{n=1}^{{N}}$, with ${\mathbf{{\hat{x}}}_{n}}=\{{\mathbf{{\hat{x}}}_{n}}[z]\}_{z=1}^{Z}$ and ${\mathbf{{\hat{x}}}_{n}}[z]\in\mathbb{R}^{D}$ are obtained by applying a set of parameterized warping functions ${\mathbf{\Omega}}={\left\{{\mathbf{\omega}_{n}}[z]:{\left\{1,\cdots,Z\right\}}\mapsto[0,1]\right\}}_{n=1}^{N}$. The Multiple Sequence Alignment (MSA) problem consists of
@@
III-B Riemannian Time Warping (RTW) Algorithm in $\mathbb{R}^{D}$
We transform ${\mathbf{X}}$ into the continuous time domain to compute the aligned signals ${\mathbf{{\hat{X}}}}$ through sinc interpolation
@@ III-B Riemannian Time Warping (RTW) Algorithm in $\mathbb{R}^{D}$
where the $\theta$-parameterized function ${\mathbf{\phi}_{\theta}}(\hat{z}):{\left[0,1\right]}\mapsto\mathbb{R}^{{N}-1}$ is modeled as a neural network, and $\hat{z}=\frac{z-1}{Z-1}$. The boundary constraint (2a) is guaranteed by design. During warping, indices of a signal may be duplicated to improve the temporal alignment, requiring a longer signal, i.e, $Z\geq{T_{\text{max}}}\,$. Hence, the continuity constraint (2b) is fullfilled with $Z={N}{T_{\text{max}}}$ [11]. To enforce the monotonicity constraint (2c), a ${\lambda}$-weighted penalty term is added to the loss as in [11]
@@
${p({\mathbf{\Omega}})}=\sum_{n=1}^{{N}}\sum_{z=1}^{Z-1}\max{\left({\mathbf{\omega}_{n}}[z]-{\mathbf{\omega}_{n}}[z+1]\,,0\right)}$ (7)
@@ IV RTW Extension for Riemannian Manifolds
Extending RTW to data on a Riemannian manifold ${\mathcal{M}}$ is not straightforward as there is no sum and scalar multiplication defined, i.e., it is not a vector space. Accordingly, the classical time warping approaches yield impractical and nonsensical results, because the warped signals and their mean would not reside on the manifold. Specifically, sinc interpolation (4), mean computation (3) and distance evaluation (9) would provide outcomes that do not respect the manifold structure. To address this, our key idea is to utilize Euclidean tangent spaces ${\mathcal{T}}$ that locally preserve distances. Then, we adjust the sinc interpolation (see IV-A), mean computation (see IV-B) and distance evaluation (see IV-C), to incorporate the curved manifold structure by operating within the tangent spaces. These novel modifications enable temporal alignment in the continuous time domain to be applied to Riemannian manifolds without affecting the computational complexity $\mathcal{O}({N}{Z})$: RTW scales linearly with ${N}$ and ${Z}$.
More specifically, a manifold ${\mathcal{M}}$ is a $D$-dimensional smooth space [14], for which there exists a tangent space ${\mathcal{T}_{{\bm{e}}}}$ for each point ${\bm{e}}\in{\mathcal{M}}$ on the manifold, with ${\bm{e}}$ being the origin of ${\mathcal{T}_{{\bm{e}}}}$. A manifold with a positive definite inner product defined on the tangent space is called Riemannian. The exponential map ${\mathbf{Exp}_{{\bm{e}}}({\bm{u}})}:{\mathcal{T}_{{\bm{e}}}}\mapsto{\mathcal{M}}$ maps a point ${\bm{u}}\in{\mathcal{T}_{{\bm{e}}}}$ from the tangent space onto the manifold. The inverse operation is the logarithmic map ${\mathbf{Log}_{{\bm{e}}}({\bm{x}})}:{\mathcal{M}}\mapsto{\mathcal{T}_{{\bm{e}}}}$. These mappings are distance-preserving, i.e. the geodesic distance between ${\bm{x}}$ and ${\bm{e}}$ is identical to ${\left\|{\mathbf{Log}_{{\bm{e}}}({\bm{x}})}\right\|}$. This feature allows indirect computation to be performed on the manifold by applying Euclidean methods in the tangent space. The formulation of the exponential and logarithmic map is manifold dependent.
@@ IV-A Windowed Sinc Interpolation on Riemannian Manifolds
The windowed sinc interpolation in (4) for a time index $z$ constitutes a weighted sum, which is not a closed operation under ${\mathcal{M}}$. Therefore, we project the original $12{\nu}+$ 1 data points from the manifold onto the tangent space ${\mathcal{T}_{{\bm{e}_{n}}[z]}}$ through ${\bm{u}_{n,z}}[m]={\mathbf{Log}_{{\bm{e}_{n}}[z]}({\bm{x}_{n}}[m])}$, where we warp and interpolate
@@
Then, the result for the time index $z$ is projected from the specific tangent space ${\mathcal{T}_{{\bm{e}_{n}}[z]}}$ back onto the manifold via
${\bm{{\hat{x}}}_{n}}[z]={\mathbf{Exp}_{{\bm{e}_{n}}[z]}({\bm{{\hat{u}}}_{n}}[z])}\,.$ (11)
Note that the choice of the tangent point ${\bm{e}_{n}}[z]\in{\mathcal{M}}$ is important to minimize undesired distortions [12]. Defining the origin of the tangent space at the center of the interpolation window as ${\bm{e}_{n}}[z]={\bm{x}_{n}}{\left[\,\lfloor{{\mathbf{\omega}_{n}}[z]\,{T_{n}}}\rfloor{}\,\right]}$, implies least distortion for points ${\bm{x}_{n}}[m]$ with the highest power of the $\operatorname{sinc}$ during interpolation. Further note that by considering a small window size $12{\nu}+$ 1 the distortion effect becomes neglectable. Obviously, the particular tangent space is reevaluated for each time index $z$ and signal $n$.
@@ IV-B Signal Mean Computation on Riemannian Manifolds
We need to adjust the mean computation of the warped signals (3), as it is not a closed operation under ${\mathcal{M}}$. Starting from an initial estimate ${\bm{{\hat{y}}}}[z]\in{\mathcal{M}}$, we utilize Gauss-Newton iterations by projecting the $z$-th data point ${\left\{{\bm{{\hat{x}}}_{n}}[z]\in{\mathcal{M}}\right\}}_{n=1}^{{N}}$ of each signal onto the tangent space ${\mathcal{T}_{{\bm{{\hat{y}}}_{n}}}}$ through ${\bm{{\hat{u}}}_{n}}[z]={\mathbf{Log}_{{\bm{{\hat{y}}}}[z]}({\bm{{\hat{x}}}_{n}}[z])}$. We obtain the center of these projections as
@@ IV-C Distance Computation on Riemannian Manifolds
The distance metric (9) is adapted as well. We utilize the geodesic distance that applies to all Riemannian manifolds
$\text{dist}_{{\mathcal{M}}}({\mathbf{{\hat{x}}}}[z],{\mathbf{{\hat{y}}}}[z])={\left\|{\mathbf{Log}_{{\bm{{\hat{y}}}}[z]}({\bm{{\hat{x}}}}[z])}\right\|}\,.$ (14)
@@
Figure LABEL:fig:overviewRTW provides a schematic overview, summarizing the proposed RTW approach. Three signals with significant temporal shifts are considered in $\mathbb{S}^{2}$, a unit sphere in $\mathbb{R}^{3}$. The color gradient from blue to red indicates the progression of time along the trajectories. We illustrate the mapping of one signal onto the tangent space for two different origins ${\bm{e}_{n}}[z]$, where the tangent space is a two-dimensional plane. As expected, distortion is noticeable only for data points far from the origin. For visualization purposes, we choose a large interpolation window ${\nu}$, indicated by the bold segments. Note that these parts are close to the origin of the tangent space but not spatially centered around it, as the window is centered in the time domain instead. Also, RTW only projects data points from a single signal onto each tangent space during interpolation, while related works often project all data points onto them, e.g., see [21, 12, 19]. Applying the warping functions ${\mathbf{\omega}_{n}}$ does not alter the shape of a signal in the tangent spaces; rather, the signal is temporally adjusted, as shown by the new color coding. RTW results in three signals on the manifold that are well aligned. After a few Gauss-Newton iterations, we obtain the signal mean, respecting the unit sphere structure.

### v3

@@ I Introduction
Time warping describes the temporal alignment of several signals – a fundamental operation for processing sequential data commonly used in speech recognition [1, 2] and human action recognition [3, 4]. In robotics, temporal alignment is crucial for learning movement primitives: it helps significantly to extract a consistent motion pattern from multiple human demonstrations that are naturally provided with temporal variations. Given multiple signals with varying lengths describing the same underlying process, each signal is warped with the goal to improve their temporal alignment [5, 6].
@@
Dynamic Time Warping (DTW) [7] finds optimal alignment through Dynamic Programming [8] but suffers from exponential complexity that prevents application to larger datasets. Alternative algorithms sacrifice global optimality to accelerate the search process by several orders of magnitude [9, 10, 11].
@@
Almost all related works focus on Euclidean data. However, many robotics problems involve temporal phenomena on Riemannian manifolds (cf. Figure LABEL:fig:overviewRTW). For example, demonstrated trajectories naturally lie on the unit sphere $\mathbb{S}^{3}$ for orientations or on symmetric positive definite matrices $\mathbb{S}^{D}_{++}$ describing a $D$-dimensional stiffness or inertia. Applying Euclidean time warping can lead to geometrically invalid interpolations (e.g., averaging quaternions in Euclidean space yields non-unit vectors) and suboptimal alignment that ignores geometric constraints [12]. While Quaternion Dynamic Time Warping (QDTW) [13] extends DTW to unit quaternions, it retains exponential complexity. To the best of our knowledge, no generic time warping method handles structured non-Euclidean data, partly because sum and scalar operations used in related works are not defined on Riemannian manifolds [14].
As our main contribution, we present Riemannian Time Warping (RTW) for efficient temporal alignment of multiple signals with varying lengths while explicitly considering the Riemannian manifold ${\mathcal{M}}$ on which the data resides. We evaluate RTW on datasets in $\mathbb{R}^{D}$, $\mathbb{S}^{D}$ and $\mathbb{S}^{D}_{++}$. RTW consistently outperforms state-of-the-art methods [10, 11] w.r.t. alignment quality and mean representation at linear complexity, and provides significant benefits for learning movement primitives. We validate our approach on real-robot experiments using a KUKA LBR iiwa manipulator in $\mathbb{R}^{3}\times\mathbb{S}^{3}$. A video is available at https://youtu.be/v_k2LwEDH9I.
@@
Notation: In the remainder, we will use square bracket notation for discrete values, e.g., ${\mathbf{x}}[t]\in\mathbb{R}^{D},{\bm{x}}[t]\in{\mathcal{M}}$, and round bracket notation for continuous functions, e.g., ${\mathbf{\phi}}(z)$.
@@ II Related Work
Most related works focus on aligning ${N}=2$ signals of length ${T}$. Despite its age, DTW [7] remains widely used due to its simplicity and effectiveness. Extending DTW to ${N}>2$ signals introduces significant computational challenges. Multiple Multi-Dimensional Dynamic Time Warping (MMDDTW) [15] computes the distance matrix over all ${N}$ dimensions. While globally optimal through Dynamic Programming, it becomes unfeasible for $N\gg 2$ due to complexity $\mathcal{O}({T}^{{N}})$. The Non-Linear Alignment and Averaging Framework (NLAAF) [16] iteratively aligns pairs of signals while adjusting previous solutions, with complexity $\mathcal{O}({N}{T}^{2})$.
Generalized Time Warping (GTW) [9], Trainable Time Warping (TTW) [10], and Neural Time Warping (NTW) [11] are newer methods that align several signals in linear time. GTW uses pre-defined basis functions, which is cumbersome for complex warpings. TTW and NTW transform discrete signals into continuous time via sinc interpolation, apply an index shift, and sample back at $Z$ uniformly distributed time points. TTW estimates discrete sine transform coefficients, while NTW uses a neural network. Both outperform GTW [10, 11].
@@ III-A Problem Statement
Consider ${N}$ discrete signals ${\mathbf{X}}=\{{\mathbf{x}_{n}}\}_{n=1}^{N}$ of lengths ${T_{n}}$, i.e. ${\mathbf{x}_{n}}=\{{\mathbf{x}_{n}}[t]\}_{t=1}^{{T_{n}}}$, where ${\mathbf{x}_{n}}[t]\in\mathbb{R}^{D}$. Signals with temporal modifications ${\mathbf{{\hat{X}}}}=\{{\mathbf{{\hat{x}}}_{n}}\}_{n=1}^{N}$ of new length ${Z}\geq{T_{\text{max}}}=\max{\left\{{T_{n}}\right\}}_{n=1}^{{N}}$, with ${\mathbf{{\hat{x}}}_{n}}=\{{\mathbf{{\hat{x}}}_{n}}[z]\}_{z=1}^{Z}$ and ${\mathbf{{\hat{x}}}_{n}}[z]\in\mathbb{R}^{D}$ are obtained by applying a set of parameterized warping functions ${\mathbf{\Omega}}={\left\{{\mathbf{\omega}_{n}}:{\left\{1,\cdots,Z\right\}}\mapsto[0,1]\right\}}_{n=1}^{N}$. The Multiple Sequence Alignment problem consists of
@@
III-B Proposed Time Warping Algorithm in $\mathbb{R}^{D}$
We transform ${\mathbf{X}}$ into the continuous time domain to compute the aligned signals ${\mathbf{{\hat{X}}}}$ through sinc interpolation [^3] Sinc interpolation provides optimal reconstruction for signals sampled at or above Nyquist rate, i.e., twice the maximum frequency of the data [21].
@@ III-B Proposed Time Warping Algorithm in $\mathbb{R}^{D}$
where the $\theta$-parameterized function ${\mathbf{\phi}_{\theta}}:{\left[0,1\right]}\mapsto\mathbb{R}^{{N}-1}$ is modeled as a neural network, and $\hat{z}=\frac{z-1}{Z-1}$. The boundary constraint (2a) is guaranteed by design. During warping, indices of a signal may be duplicated to improve the temporal alignment, requiring a longer signal, i.e., $Z\geq{T_{\text{max}}}\,$. Hence, the continuity constraint (2b) is fulfilled with $Z={N}{T_{\text{max}}}$ [11]. To enforce the monotonicity constraint (2c), a ${\lambda}$-weighted penalty term is added to the loss as in [11]
@@
${p({\mathbf{\Omega}})}=\sum_{n=1}^{{N}}\sum_{z=1}^{Z-1}\max{\left({\mathbf{\omega}_{n}}[z]-{\mathbf{\omega}_{n}}[z+1]\,,0\right)}\,.$ (7)
@@ IV RTW Extension for Riemannian Manifolds
Extending RTW to data on a Riemannian manifold ${\mathcal{M}}$ is not straightforward as there is no sum and scalar multiplication defined, i.e., it is not a vector space. Accordingly, the classical time warping approaches yield geometrically invalid results, because the warped signals and their mean would not reside on the manifold. Specifically, sinc interpolation (4), mean computation (3) and distance evaluation (9) would provide outcomes that do not respect the manifold structure. To address this, our key idea is to utilize Euclidean tangent spaces ${\mathcal{T}}$ that locally preserve distances. Then, we adjust the sinc interpolation (see IV-A), mean computation (see IV-B) and distance evaluation (see IV-C), to incorporate the curved manifold structure by operating within the tangent spaces. These novel modifications enable temporal alignment in the continuous time domain to be applied on Riemannian manifolds without affecting the computational complexity $\mathcal{O}({N}{Z})$: RTW scales linearly with ${N}$ and ${Z}$.
More specifically, a manifold ${\mathcal{M}}$ is a $D$-dimensional smooth space [14], for which there exists a tangent space ${\mathcal{T}_{{\bm{p}}}}$ for each point ${\bm{p}}\in{\mathcal{M}}$ on the manifold, with ${\bm{p}}$ being the origin of ${\mathcal{T}_{{\bm{p}}}}$. A manifold with a positive definite inner product defined on the tangent space is called Riemannian. The exponential map ${\mathbf{Exp}_{{\bm{p}}}}:{\mathcal{T}_{{\bm{p}}}}\mapsto{\mathcal{M}}$ maps a point from the tangent space onto the manifold. The inverse operation is the logarithmic map ${\mathbf{Log}_{{\bm{p}}}}:{\mathcal{M}}\mapsto{\mathcal{T}_{{\bm{p}}}}$. These mappings are distance-preserving, i.e., the geodesic distance between the point ${\bm{p}}$ and any other point is maintained. This feature allows indirect computation to be performed on the manifold by applying Euclidean operations in the tangent space. The formulation of the exponential and logarithmic map is manifold dependent.
@@ IV-A Windowed Sinc Interpolation on Riemannian Manifolds
The windowed sinc interpolation in (4) for a time index $z$ constitutes a weighted sum, which is not a closed operation under ${\mathcal{M}}$. Therefore, we project the original $12{\nu}+$ 1 data points from the manifold onto the tangent space ${\mathcal{T}_{{\bm{p}_{n}}[z]}}$ through ${\bm{u}_{n,z}}[m]={\mathbf{Log}_{{\bm{p}_{n}}[z]}({\bm{x}_{n}}[m])}$, where we warp and interpolate
@@
Then, the result for the time index $z$ is projected from the specific tangent space ${\mathcal{T}_{{\bm{p}_{n}}[z]}}$ back onto the manifold via
${\bm{{\hat{x}}}_{n}}[z]={\mathbf{Exp}_{{\bm{p}_{n}}[z]}({\bm{{\hat{u}}}_{n}}[z])}\,.$ (11)
Note that the choice of the tangent point ${\bm{p}_{n}}[z]\in{\mathcal{M}}$ is important to minimize undesired distortions [22]. Therefore, we initialize the origin of the tangent space at the center of the interpolation window as ${\bm{p}_{n}}[z]={\bm{x}_{n}}{\left[\,\lfloor{{\mathbf{\omega}_{n}}[z]\,{T_{n}}}\rfloor{}\,\right]}$, and refine it through few Gauss-Newton iterations ${\bm{p}_{n}}[z]\leftarrow{\bm{{\hat{x}}}_{n}}[z]$ to approximate the optimal tangent point that minimizes distortions, i.e., we compute a weighted Fréchet mean within the interpolation window. Due to the continuity constraint of the warping functions ${\mathbf{\omega}_{n}}$, the Fréchet mean is constrained to lie on the sinc interpolated line connecting the three adjacent discrete data points ${\bm{x}_{n}}[\lfloor{{\mathbf{\omega}_{n}}[z]{T_{n}}}\rfloor{}-1]$, ${\bm{x}_{n}}[\lfloor{{\mathbf{\omega}_{n}}[z]{T_{n}}}\rfloor{}]$ and ${\bm{x}_{n}}[\lfloor{{\mathbf{\omega}_{n}}[z]{T_{n}}}\rfloor{}+1]$. Further note that the optimal tangent space is reevaluated for each time index $z$ and signal $n$, instead of projecting all data points onto one single tangent space.
@@ IV-B Signal Mean Computation on Riemannian Manifolds
Similar to the sinc interpolation, the mean computation (3) is adjusted by utilizing Gauss-Newton iterations. Starting from an initial estimate ${\bm{{\hat{y}}}}[z]\in{\mathcal{M}}$, e.g., ${\bm{{\hat{y}}}}[z]={\bm{{\hat{x}}}_{1}}[z]$, the $z$-th data point ${\left\{{\bm{{\hat{x}}}_{n}}[z]\in{\mathcal{M}}\right\}}_{n=1}^{{N}}$ of each signal is projected onto the tangent space ${\mathcal{T}_{{\bm{{\hat{y}}}_{n}}}}$ through ${\bm{{\hat{u}}}_{n}}[z]={\mathbf{Log}_{{\bm{{\hat{y}}}}[z]}({\bm{{\hat{x}}}_{n}}[z])}$. We obtain the center of these projections as
@@ IV-C Distance Computation on Riemannian Manifolds
The distance metric (9) is adapted as well. We utilize the geodesic distance on Riemannian manifolds
$\text{dist}_{{\mathcal{M}}}({\bm{{\hat{x}}}}[z],{\bm{{\hat{y}}}}[z])={\left\|{\mathbf{Log}_{{\bm{{\hat{y}}}}[z]}({\bm{{\hat{x}}}}[z])}\right\|}_{{\bm{{\hat{y}}}}[z]}\,,$ (14)
where the norm is computed with respect to the Riemannian metric at ${\bm{{\hat{y}}}}[z]$, which depends on the manifold’s geometry.
@@
Figure LABEL:fig:overviewRTW provides a schematic overview, summarizing the proposed RTW approach. Three similar signals with significant temporal shifts are considered in $\mathbb{S}^{2}$, a unit sphere in $\mathbb{R}^{3}$. The color gradient from blue to red indicates the progression of time along the trajectories. We illustrate the mapping of one signal onto the tangent space for two different origins ${\bm{p}_{n}}[z]$, where the tangent space is a two-dimensional plane. As expected, distortion is noticeable only for data points far from the origin. For visualization purposes, we choose a large interpolation window ${\nu}$, indicated by the bold segments. Note that these parts are close to the origin of the tangent space but not spatially centered around it, as the window is centered in the time domain instead. Furthermore, RTW only projects data points from a single signal onto each tangent space during interpolation, while related works often project all data points onto them, e.g., see [23, 12, 19]. Applying warping functions ${\mathbf{\omega}_{n}}$ does not alter the shape of a signal in the tangent spaces; rather, signals are temporally adjusted, as shown by the new color coding. RTW results in three signals on the manifold that are well aligned. After a few Gauss-Newton iterations, we obtain the signal mean, respecting the unit sphere structure.